OpenStudy (anonymous):

factor the trinomial completely. factor the trinomial completely. @Mathematics

OpenStudy (anonymous):

problem on attachment. please help

OpenStudy (anonymous):

I got 2(n-3)^2 but looks like that is wrong

OpenStudy (anonymous):

its 3(n-2)^2

OpenStudy (anonymous):

ooh opps!

OpenStudy (agreene):

your coefficient is incorrect. You should arrive to \[3(n-2)^2\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

3n^2-12n+12 =3(n^2 -4n+4) =3(n-2)(n-2) =3(n-2)^2 ans